In a study about the hourly wage (X, in rands) paid to young people aged 14-18 with a holiday job,
the following summarised results were measured for a random sample of 400 young people.
X
400
iD1 xi D 3 296 X
400
iD1 x2i D 27 833:35
Two years ago, the mean hourly wage paid to young people with a holiday job was R8:05, with
standard deviation R1:25. The question arises whether the current hourly wage  is greater than
R8:05. In the questions below, assume that the variation in the hourly wages has not changed.Calculate the mean and standard deviation of the data.
(b) Use a 95% confidence interval to answer the question.
(c) Perform a hypothesis test at the 5% level of significance to answer the question (Make sure
you indicate all steps).
(d) Do you need to know whether X is normally distributed? Why/why not?
1
Expert's answer
2020-03-30T10:33:02-0400

Given that


"\sum_{\mathclap{1\le i\le 400}} X_i=3296, \ \ \ \ \sum_{\mathclap{1\le i\le 400}} X_i^2=27833.35"

a) calculate the mean and standard deviation of the data


"\mu=E(X)={1 \over n}\sum_{\mathclap{1\le i\le n}} X_i={1 \over 400}\sum_{\mathclap{1\le i\le 400}} X_i="

"={3296 \over 400}=8.24"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2="


"={27833.35 \over 400}-(8.24)^2=1.685775"

"\sigma=\sqrt{\sigma^2}=\sqrt{1.685775}\approx1.2984"

b)We need to construct the "95\%" confidence interval for the population mean "\mu" with unknown population variance.

The critical value for "\alpha=0.05" and "df=n-1=399" degrees of freedom is


"t_c=t_{1-\alpha/2;n}=1.966"

The corresponding confidence interval is computed as shown below:


"CI=\big(\bar{X}-t_c\times{s \over \sqrt{n}},\ \bar{X}+t_c\times{s \over \sqrt{n}}\big)="

"=\big(8.24-1.966\times{1.2984 \over \sqrt{400}},\ 8.24+1.966\times{1.2984 \over \sqrt{400}}\big)="

"=(8.112,\ 8.368)"

If we use z-test.

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96."

The corresponding confidence interval is computed as shown below:


"CI=\big(\bar{X}-z_c\times{\sigma \over \sqrt{n}},\ \bar{X}+z_c\times{\sigma \over \sqrt{n}}\big)="

"=\big(8.24-1.96\times{1.2984 \over \sqrt{400}},\ 8.24+1.96\times{1.2984 \over \sqrt{400}}\big)="

"=(8.113,\ 8.367)"

c) The following null and alternative hypotheses need to be tested:

"H_0:\mu\leq8.02"

"H_1:\mu>8.02"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation "\sigma=1.25" will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a right-tailed test is "z_c=1.64."

The rejection region for this right-tailed test is "R=\{z:z>1.64\}"

The z-statistic is computed as follows:


"z={\bar{X}-\mu \over \sigma/\sqrt{n}}={8.24-8.05 \over 1.25/\sqrt{400}}=3.04"

Since it is observed that "z=3.04>1.64=z_c," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is greater than "8.05," at the "0.05" significance level.

Using the P-value approach: The p-value for "z=3.04" is "p=0.001183," and since "p=0.001182<0.05," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is greater than "8.05," at the "0.05" significance level.

The 95% confidence interval is


"CI=\big(\bar{X}-z_c\times{\sigma \over \sqrt{n}},\ \bar{X}+z_c\times{\sigma \over \sqrt{n}}\big)="

"=\big(8.24-1.96\times{1.25 \over \sqrt{400}},\ 8.24+1.96\times{1.25 \over \sqrt{400}}\big)="

"=(8.1175,\ 8.3625)"


d) When the sample size is large, the "z-"tests are easily modified to yield valid test procedures without requiring either a normal population distribution or known standard deviation.